Considering the following financial information for Atlas Awesome Manufacturing, Inc. and Delilah Superior Manufacturing Inc. Both companies are in the same industry and have identical operating income of $8.4 million. Atlas finances its $15 million in assets with $2 million debt ( on which it pays 9 percent interest) and 13 million in equity. Delilah finances its $15 million in assets with $12 million in debt ( on which it pays 8 percent interest). Both companies pay 32 percent tax on their taxable income.

Calculate the following:

1. Each firms net income

2. The income each firm has available to pay its debt holders and stock holders (the firms asset funders)

3. The returns available to the asset funders on their investment in each company (the return on asset-funders investment)

4. Which company offers a higher return on investment to its asset funders? Explain why this company is able to offer a higher rate of return on investment to its asset funders.

Please show all calculations in digital typed format. Please don't use hand writing. Thank you so much.
Answer #1

A) Let us consider Atlas first

Operating Income EBIT= $ 8.4 Million=$8400,000

Interest=0.09*2 Million= $180,000

EBT= $840,000- $ 180,000= $8,220,000

Tax=0.32*8,220,000=$2,6,46,840

Net income=8,220,000-2,646,840=$5,173,160

Delliah:

Operating Income EBIT= $ 8.4 Million=$8400,000

Interest =$ 12 Million *0.08=960,000

EBT=$7,440,000

Tax=0.32*7440,000=2,380,000

Net Income= 5,060,1000